A plane can fly 2486 miles in the same time as it takes a car to go 525 miles.If the car travels 280 mph slower than the plane, find the speed of the plane.
Answer by jorel1380(3719)

525/x=2486/x+280
525x+147000=2486x
1961x=147000
x=74.96
x+280=354.96..